OpenStudy (anonymous):

ABC is a right triangle. if AC =4 and BC=10,find AB leave the answer in simplest radical form a116 b4 c2 racial 21 d2 racial 26

OpenStudy (anonymous):

c

OpenStudy (anonymous):

may you explain it to me plz

OpenStudy (anonymous):

a^2 + 100 = 16

OpenStudy (anonymous):

damn i got sqrt of -84....

OpenStudy (anonymous):

hmmm...well i somehoe got sqrt of 84 then i jus did 84/2 = 42 then i did 42/2 = 21 you have two 2's there and a left over 21 so you do 2 sqrt of 21

OpenStudy (anonymous):

thank once again :))

OpenStudy (anonymous):

np